OpenStudy (anonymous):

A pair of dice is thrown. What is the probability that the first die is even and the second is 5? Wouldn't it be 1/36? Or do I have to simplify it more?

hartnn (hartnn):

NO, its not 1/36 p(even) = ? its not 1/6

OpenStudy (anonymous):

3/6? = P(even)

hartnn (hartnn):

yes, and 3/6=1/2 now required probability = 1/2 * 1/6 = ??

OpenStudy (anonymous):

1/12?

hartnn (hartnn):

yup.

OpenStudy (anonymous):

Does it need to be simplified or anything?

hartnn (hartnn):

nopes. why 1/12 isn't in options?

OpenStudy (anonymous):

It is.